TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint4x2007x06x27x18x51

going on activities in sector 45

Going on activities in sector 45

  • DFBAH and MB (diode)
    • checked and closed by Wednesday 27th;
    • Leak test on Thursday 28th;
    • Purging on Friday 29th evening and all the week end long;
    • Refilling on Monday 2nd July;
    • On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB);
  • CRYO-instrumentation
    • Initialization of RF valves on Wednesday 27th or Thursday 28th;
    • Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
    • Commissioning;
    • Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
  • TS/EL intervention
    • Changing of the connectors of a BPM (LSS4R);
    • Placing of some Radmon cables on Wednesday 27th, Thursday 28th and Friday 29th (LSS5L);
    • Connection check of the instrumentation cables of BCT;
  • TIM camera
    • General check on Monday 2nd July;
    • Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
  • Safety Inspection
    • Monday 2nd July at 2pm;
  • RAMSES
    • Check of the equipment on Tuesday 26th and Wednesday 27th.
  • CV Intervention
    • Thursday 28th June from 2pm to 4pm
  • QUI
    • Caps removal, sleeve installation and then closure of the QUI performed on Thursday 28th.
    • The new mechanism for the header B valve is being prepared and should be ready for Friday 29th or early next week.

Comments:

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2007-06-28 - MatteoSolfaroli
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login